OpenStudy (anonymous):

If x + 3 / 3 = y + 2 / 2 then x/3 = ??? These are set as proportions please help me i’ve been stuck for hours :(

OpenStudy (campbell_st):

well split the 1st fraction (x +3)/3 = x/3 + 3/3 = x/3 + 1 now look at the right... same thing (y + 2)/2 = y/2 + 2/2 = y/2 + 1 now simplify it.

OpenStudy (anonymous):

Would i simplify it as if i'm solving a proportion?

OpenStudy (campbell_st):

well the problem becomes x/3 + 1 = y/2 + 1 subtract 1 from both sides of the equation and you have a solution

OpenStudy (anonymous):

Thank you so much!